The speed of sound as measured by a student in the laboratory on a winter day is 340 m s-1 when the room temperature is 17°C. What speed will be measured by another student repeating the experiment on a day when the room temperature is 32°C?

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. V base 1 = 330 m/s, V base 2 = ? T base 1 = 273 + 17 = 290 K, T base 2 = 272 + 32 = 305 K We know v ∝ √T √V abse1/√V base 2 = √T abse1/√T base 2⇒ V base 1 * √T base 2/√T base 1 = 340 *√305/290 = 349 m/s.